MODS1 Dark Commissioning and Spring SV Run: UTC 2011 May 29

Observer(s): R. Pogge (OSU), M. Wagner (OSU/LBTO), M. Pedani (LBTO)
Telescope Operator: G. Bechetti (LBTO)
Telescope Support: M. Midkiff (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

Night 1 of 4 devoted to dark-time commissioning and early science verification observations for the LBTO community.

Night lost to a combination of high winds and failure of the LDG rotator. See below for the latter.

Preparations

Replaced the commissioning long-slit masks (LS5x60x0.3, 0.6, and 1.2) with the new "facility" long-slit masks (LS5x60x0.3, 0.6, 0.8, 1.0, and 1.2 arcsec), removed 3 test masks and inserted 3 science MOS masks for the spring SV program. A 4th mask was set aside because it was clearly made wrong (something funny about the Gerber file, not a problem at URIC cutting the mask).

Checked and re-seated the Hg(Ar) calibration lamp which had slipped out of its mount on the integrating sphere. All other pen-ray type calibration lamps were inspected and found to be firmly seated.

Trained available support staff (Doug and Marco) in the mask exchange procedure for MODS. The process went very smoothly, and it will be written up. Some problems getting a laptop to connect via DHCP on Level 5 (which we've been doing without problems all this past year). All 3 live ethernet cables we found would not negotiate the DCHP connection up there. We finally had to hardwire ourselves into MODS and use a fixed IP (good thing I had the MODS lab laptop with us). IssueTrak 3410 was submitted on this problem.

The LDG rotator started showing a number of faults. Dan Cox was called in and it is behaving badly. Wind is at or over limits at sunset, so we're not observing anyway, but still...

Details

All times are UTC unless otherwise noted.

0130 - During startup, we immediately got LDG rotator faults - no problems earlier today doing a few small rotations at zenith. After much trying of usual solutions, Dan Cox was called in to help investigate. They've started working on it.

0330 - Winds are sustained around 20m/s with gusts as high as 25m/s during the past hour. Definitely picking after sunset. We are still closed by winds and by a serious rotator problem being worked by folks downtown as well as up here.

0430 - Still working the rotator problem. Wind gusting above limits (25m/s). Of course it is clear tonight...

0655 - The rotator problem is the drive module. Mike just called the night. There is a spare module on the mountain that will be installed tomorrow and checked out. The winds gusting above limits take some of the edge off, but it is clear outside. Should the winds drop, there is nothing we can do.

-- RichardPogge - 28 May 2011
Topic revision: r3 - 29 May 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback